6. Find the least squares line associated with the following data. Show all work for any credit. X 12 2.1 3.2 6.3
Answer #1

6

x y xy x2
12 2.1 25.2 144
10 3.2 32 100
6 6.3 37.8 36
2 9.1 18.2 4

To find the least squares line, the following three formulas are required

m= [n(\sumxy)-(\sumx)(\sumy)]/ [n(\sumx2)-(\sumx)2

b= (\sumy)-m(\sumx)/n

y=mx+b(formula for equation of a line: slope-intercept form)

m: slope of line

b: y-intercept

n=4

m=[4(113.2)-30(20.7)]/[4(284)-(30)2]

=- 168.2/236=0.7127

Substituting the value of m in the equation for b we get

b=[ 20.7-(-168.2/236*30)]/4

=10.520

Substituting the above values in the equation y=mx+b

y= 0.7127x+10.520 is the least squares line associated with the given data